OpenStudy (anonymous):

Hi Guys. Solve for x. 2logx=log2+log(6x)-10) Need a lower and upper value....drawing a blank on where to begin.

OpenStudy (anonymous):

i can't really read it is it \[2\log(x)=\log(2)+\log(6x-10)\]?

OpenStudy (anonymous):

2 log x = log 2 + log(6x − 10) need an upper and lower value. I thought to start I should move the 2 log x to the other side but I don't know what then.

OpenStudy (anonymous):

that is probably the best idea, get \[2\log(x)-\log(6x-10)=\log(2)\] then combine the stuff on the left in to a single log using \[n\log(x)=\log(x^n)\] and \[\log(A)-\log(B)=\log\left(\frac{A}{B}\right)\]

OpenStudy (anonymous):

i'll try it. one moment

OpenStudy (anonymous):

right on the money....it became x^2-12x+20....values on 2 and 20. I may have one or two more I could use a bit of direction on. mind if I look u up later? thanks

OpenStudy (anonymous):

ok.....log2(x + 8) − log2(x − 8) = 3 initially I am thinking make everything base 10 to get 2(x+8)-2(x-8)=3. Am I on the right track?

OpenStudy (anonymous):

leave it as base 2

OpenStudy (anonymous):

combine in a single log like before

OpenStudy (anonymous):

log 2 ((x+8)/(x-8))=3?